### v1

# Iterative Approach to Reconstructing Neural Disparity Fields from Light-field Data

Abstract
This study proposes a neural disparity field (NDF) that establishes an implicit, continuous representation of scene disparity based on a neural field and an iterative approach to address the inverse problem of NDF reconstruction from light-field data. NDF enables seamless and precise characterization of disparity variations in three-dimensional scenes and can discretize disparity at any arbitrary resolution, overcoming the limitations of traditional disparity maps that are prone to sampling errors and interpolation inaccuracies. The proposed NDF network architecture utilizes hash encoding combined with multilayer perceptrons to capture detailed disparities in texture levels, thereby enhancing its ability to represent the geometric information of complex scenes. By leveraging the spatial-angular consistency inherent in light- field data, a differentiable forward model to generate a central view image from the light-field data is developed. Based on the forward model, an optimization scheme for the inverse problem of NDF reconstruction using differentiable propagation operators is established. Furthermore, an iterative solution method is adopted to reconstruct the NDF in the optimization scheme, which does not require training datasets and applies to light-field data captured by various acquisition methods. Experimental results demonstrate that high-quality NDF can be reconstructed from light-field data using the proposed method. High-resolution disparity can be effectively recovered by NDF, demonstrating its capability for the implicit, continuous representation of scene disparities.
Index Terms:
Neural Disparity Field, Disparity estimation, Light Field
I Introduction
Computational imaging innovatively incorporates computer vision and signal processing techniques into imaging systems to reconstruct high-dimensional optical signals that carry scene information through novel imaging mechanisms, optical paths, and reconstruction methods, thereby achieving efficient and high-precision observation of scenes [1]. Computational imaging surpasses the limitations of classical imaging in terms of signal dimensions, scale, and resolution.
From the perspective of imaging methodology, light-field imaging operates as a close-proximity passive direct line-of-sight imaging technology that relies on ambient light to capture scene content within its visual ambit. The characteristic features included narrower baselines, increased uniformity in viewpoint sampling, and diminished occlusions. Notably, substantial domain discrepancies emerge in the light-field data generated by different cameras owing to inherent variations in their imaging modalities, structural designs, and internal parameters.
Deep-learning-based reconstruction methods rely on labeled training data and lack exploration of the inherent consistency within light-field data. Consequently, networks trained using data from one camera model may not be generalizable to other camera models. Given that individual light-field data contain abundant information, exploiting the full potential of their spatial-angular consistency empowers a network to reduce its reliance on specific camera models, thereby promoting generalization across different camera models.
The neural field [2] represents the scene to be reconstructed as a continuous implicit function, achieving advanced visual quality by the neural network parameterization. The continuous implicit function representation ensures continuity and is not limited by resolution constraints. By employing neural network parameterization of signal values, this representation method adopts a ”discrete-to-continuous” computational approach and utilizes differentiability for solving inverse problems. Compared with traditional methods, the neural field exhibits superior precision, lower complexity, and increased robustness. Furthermore, compared with deep learning approaches, their advantages include not requiring extensive training datasets, being unrestricted by data formats, having fewer parameters, and exhibiting high robustness.
Therefore, we utilize a fully connected neural network to parameterize the disparity. The network constructs a neural disparity field (NDF) by incorporating hash encoding and multilayer perceptrons (MLPs) with activation functions. This approach implicitly represents the disparity contained in the light field and exploits the spatial-angular consistency within it. The inverse problem-solving model based on the NDF was established to reconstruct the disparity from the light field. This model was trained and iteratively solved through differentiability, effectively transforming the disparity estimation problem into an optimization problem for the NDF. The framework for the implicit representation of an NDF is shown in Fig.1.
Figure 1: Framework for the implicit representation of neural disparity fields. Red ray bundles indicate the propagation of disparities from the visible views towards the central view.
By exploiting the spatial-angular consistency within individual light-field data, the proposed method demonstrates exceptional performance in reconstructing the NDF. This reconstructed field can restore high-quality explicit discrete disparity maps. Unlike depth-learning-based methods for disparity estimation, the NDF reconstruction method does not depend on training datasets. It can be applied to two-plane parameterized light-field data using any collection method. By utilizing the implicit representation of the neural field, this disparity reconstruction method departs from the traditional approach of representing disparity as discrete points, overcoming the inevitable sampling, interpolation, and other discretization calculation issues of explicit discrete disparity and establishing a continuous disparity surface with implicit expression. The innovations of this study are as follows:
  1. 1.
    Disparity is parameterized using neural networks to establish an NDF, that creates an implicit continuous representation of scene disparity functions. This NDF enhances the expressive capability of complex details of the disparity surface and effectively reduces computational errors introduced by discretization. Unlike traditional hierarchical disparity search estimation methods, a representation based on the NDF allows the formation of a differentiable propagation forward problem model, inverse problem optimization modeling, and iterative solving of the reconstruction algorithm for continuous implicit disparity functions.
  2. 2.
    From the perspective of spatial-angular consistency, a forward problem model for the disparity propagation of the NDF under non-ideal scene conditions (with occlusion and noise) is constructed. This forward-problem model, described by differentiable propagation operators, can propagate any viewpoint image to a central view via the NDF.
  3. 3.
    Based on the forward problem model, an optimization model for the inverse problem of the NDF was established using differentiable propagation operators. An iterative approach was proposed to solve the disparity reconstruction from the light-field. This method offers a novel solution to the disparity reconstruction problem. The iterative reconstruction method based on the inverse problem does not rely on training datasets and can be applied to two-plane parameterized light-field data obtained by any collection method. Furthermore, it allows the application of inverse problem-solving strategies in disparity reconstruction.
II Related work
The 7D plenoptic function[3] was initially used to characterize the light fields. Over time, this function has been transformed into a two-plane four-parameter light field model[4], with the majority of commonly used methods relying on four-dimensional(4D) two-plane light fields[5]. The utilization of light field technology is crucial for providing depth information. To ensure generality, the proposed method is demonstrated using the central viewpoint depth and possesses the ability to reconstruct depth information from alternative viewpoints. Additionally, for light fields, there exists a one-to-one correspondence between the scene depth and disparity, which accurately represents the three-dimensional(3D) spatial structure of the scene. Existing methods can be broadly classified into two categories: rule-based traditional methods and deep-learning-based methods.
II-A Rule-based traditional methods
Disparity estimation algorithms that are rule-based traditional methods adhere to a computational approach known as ”discrete-to-discrete”. They aimed to estimate the corresponding discrete disparity images from discretized light-field data. However, this computational approach is influenced by the resolution limitations of disparity and unavoidably introduces various errors. A common source of error is sub-pixel interpolation, given that the accuracy of interpolation influences the estimation precision of the algorithm. For instance, Jeon et al. [6]proposed the use of the Fourier phase-shifting theory to replace bilinear interpolation, resulting in improved estimation accuracy. Currently, researchers have not noted advancements in the mitigation of discretization errors.
Early disparity estimation followed rule-based estimation methods, exploring the structural information of light-field data to assess the consistency among different views. Wanner et al. [7] introduced structural tensors into depth estimation from light-field data, enhancing the global consistency of the reconstructed depth by constraining the continuity of depth from different subview images of the light field. However, owing to the insufficient utilization of angular information in light-field data, the estimation performance of this algorithm requires further improvement. Heber et al. [8] introduced principal component analysis for the computation of stereo matching. They focused on subview images that impacted matching errors while disregarding those with minor influences. This approach improves the robustness of the algorithm and to some extent, its precision for stereo matching. To overcome the challenges of matching small baselines between subview images in light fields, Jeon et al. [6] applied Fourier phase-shifting theory to improve the accuracy of the matching computation. They adaptively aggregated the gradient costs based on the angular coordinates of the light fields and incorporated feature correspondences between subaperture images as additional constraints. Tao et al. [9]examined the properties of two depth cues derived from the correspondence and defocus in light-field data. Their findings demonstrate that these cues yield complementary depth information. Consequently, researchers have suggested merging the two depth cues to incorporate both angular and spatial information from the light field, thereby enhancing the performance of the algorithm for disparity estimation. However, this algorithm requires manual parameter adjustment to blend different results, thereby affecting its robustness across different scenes. In addition, it does not account for occlusion effects, leading to estimation errors in occluded regions.
To mitigate the errors arising from occlusions, Wang et al. [10] expanded the work of [9] by developing a model for the occlusion effect in light fields. They proposed an anti-occlusion algorithm by studying the relative positional relationships between different subview images in a light field and incorporating prior knowledge of object occlusion into stereo-matching algorithms. Although this algorithm partially addresses the challenges associated with occlusion in depth estimation by modeling occlusion effects in light fields, its straightforward occlusion detection criteria result in limited robustness across different scenes. To further mitigate the impact of occlusion on depth estimation performance, Sheng et al. [11] devised an estimation algorithm based on epipolar plane images (EPIs) combined with the local consistency of scene depth to address occlusion issues. Zhang et al. [12] considered the structure of EPIs and proposed a local depth estimation method based on a spinning parallelogram operator(SPO). Utilizing the regions segmented by the SPO in EPIs, they located lines representing depth information by maximizing the distribution distances between areas. Liu et al. [13] proposed an iterative scheme for scene depth reconstruction based on a 4D light field. The objective function of their optimization problem consisted of a fidelity term representing matching, along with penalty terms comprising gradient and classification terms. It iteratively corrects mismatched pixels to minimize the objective function, ultimately achieving accurate depth estimation. Subsequently, Liu et al. [14] mapped the pixel differences between matching regions to radial basis functions, where pixels with larger differences in occluded and noisy areas were mapped to smaller values in the radial basis functions, thereby effectively achieving anti-occlusion resistance and antinoise.
Although rule-based traditional disparity estimation algorithms achieve relatively high precision in depth estimation by exploiting the interrelationships among sub-aperture images, optimizing stereo matching computations, and addressing occlusion effects, their performance is still limited owing to the insufficient exploitation of angular and spatial information within the light field. Additionally, these methods, which use a ”discrete-to-discrete” approach, fail to overcome the resolution limitations inherent in the current representation of disparity and remain vulnerable to errors caused by data discretization. By contrast, neural fields with implicit representations, characterized by their inherent continuity, overcome the limitations imposed by finite-resolution discrete disparities and offer enhanced adaptability to complex variations in disparity and precision in their depiction.
II-B Deep learning-based methods
With the advancements made in deep learning techniques for diverse computer vision tasks, there has been increasing interest in depth estimation from light-field data. Deep-learning-based algorithms primarily depend on supervised learning approaches, which require a large amount of training data annotated with accurate depth labels.
Heber et al. [15] introduced convolutional neural networks (CNNs) to address the task of depth estimation from light field data. Their approach utilizes supervised learning by providing a network with subaperture images in the horizontal and vertical directions and the corresponding depth labels. Data-augmentation techniques involving rotation, translation, brightness, and color adjustments have been proposed to improve the training process. Although the algorithm demonstrated superior performance on synthetic datasets compared to traditional algorithms, it required a large amount of depth-labeled training data, resulting in substantial data generation costs. In additionally, the domain gap between synthetic and real data poses challenges to the applicability of the algorithm in practical imaging systems. Subsequently, Shin et al. [16] proposed a multi-branch network structure with four branches dedicated to extracting features from light-field images in horizontal, vertical, and diagonally tilted directions. These features were then combined using CNNs to reconstruct the scene depth. Although the method showed improved estimation performance on synthetic data compared with traditional disparity estimation algorithms, it exhibited limitations in occluded and specular regions.
Alperovich et al. [17] developed a fully convolutional autoencoder for simultaneous depth and reflectance estimations. Although this supervised approach partially addresses the occlusion effects, its generalization ability is limited. Peng et al. [18] introduced unsupervised learning for depth estimation using light-field data. They proposed a combined loss function that imposed compliance and divergence constraints on the disparity between the warped sub-aperture images and the central view. This approach alleviates the challenge of acquiring labeled data and improves the generalizability of the algorithm across different imaging systems. Zhou et al. [19] utilized geometrically constrained EPIs to train networks and reconstruct depth information without requiring ground-truth depth labels. Li et al. [20] introduced an Oriented Relation Module that captured spatial relationships in vertical and horizontal EPIs. They enhanced training effectiveness and depth-estimation performance using refocusing-based data augmentation. Tsai et al. [21] presented a viewpoint-selection module that generated attention maps to determine the importance of each view for accurate depth estimation. This attention-driven approach improves the effective utilization of light field data. Li et al. [22] developed a lightweight model to estimate the disparity. The model employed physics-based multiscale cost-volume convolutional aggregation and an edge-guided subnetwork to enhance the geometric detail recovery near the edges and the overall performance. Huang et al. [23] designed an end-to-end trainable lightweight convolutional network that integrated cost volumes and attention modules. The designed network architecture enables the inference of light-field inputs at varying angles and facilitates detailed recovery in occluded regions.
The acquisition of the training and labeling data required by existing deep-learning algorithms poses a challenge, thus limiting their precision and applicability. Most deep learning-based approaches depend on data sourced exclusively from particular camera models during training, impeding their capacity to acclimate to various camera models and diminishing their generalization ability.
II-C Neural fields
The neural radiance field [24] represents the scene to be reconstructed as a continuous implicit function parameterized by a neural network, achieving advanced visual quality and inspiring numerous subsequent works [25, 26, 27, 28, 29, 30]. Among these, mip-NeRF [25] reduces the multiscale representation of image aliasing in the neural radiance field. Mip-NeRF 360 [26] extended mip-NeRF to address poor performance in unbounded scene representations. Ref-NeRF [27] addresses specular and reflection issues in novel view synthesis. The instant-NGP [28] model enhances the training speed of neural radiance field models. TensorRF [29] decomposes the radiance field tensor into multiple lower-order tensor components to obtain a more compact scene representation. Factor Fields [30] is a unified framework based on coordinate-input networks, allowing the creation of powerful new signal representations and achieving improvements in approximation quality, compactness, and training time.
Continuous implicit functions parameterized by neural fields exhibit continuity and are not subject to limitations imposed by finite resolution, showing advantages over discrete representations (e.g., pixel color [31, 32], light field [33, 34, 35], signed distance [36], and occupancy [37]) in certain applications[38]. The core idea involves training MLPs with activation functions to map the input coordinates to the desired target values. Similar to existing coordinate-based representation methods, our method learns to map the coordinates of the central view to the corresponding disparities within the central view scene.
III Method
Traditional methods for disparity estimation from light field data follow a ”discrete-to-discrete” approach, which fails to overcome the resolution limitations inherent in disparity representation. By contrast, with their continuous nature, neural field implicit representations provide superior expression capabilities compared to traditional discrete disparity representations, as they are not bound by finite resolution constraints, thus enabling robust modeling of complex disparity variations.
III-A Neural disparity field representation
In contrast to traditional implicit representations of 3D shapes, disparity refers to a two-dimensional (2D) geometric manifold in 3D space that is associated with the depth of the scene. It describes the differences between viewpoints instead of directly representing 3D geometric shapes. The implicit continuous parameterization of NDFs allows the modeling of continuous disparities, making it more appropriate for describing continuous variations in disparity within light fields compared to the conventional 2D matrix representations.
The disparity inherent in a scene within a light field is parameterized as a continuous 2D implicit function using neural fields, termed NDFs. This is achieved by approximating the continuous 2D implicit function using a multiresolution grid and the MLPs network $\mathbf{F}_{\mathbf{\Theta}}$ with activation functions, defined as
$\mathbf{F}_{\mathbf{\Theta}}:\mathbf{x}\rightarrow{d(\mathbf{x})},$ (1)
where the network takes 2D coordinate positions $\mathbf{x}=(x,y)$ as input and outputs the disparity ${d(\mathbf{x})}$ at that point. By optimizing the network weights $\mathbf{\Theta}$, this representation maps each input 2D coordinate $\mathbf{x}$ to its corresponding disparity ${d(\mathbf{x})}$.
III-B Inverse problem solving for disparity reconstruction
In the light field, disparity information is shared across various subviews, allowing viewpoint transitions through disparity propagation. In the two-plane four-parameter light field, assuming that occlusion and camera noise are disregarded, the relationship between image $L_{(u_{0},v_{0})}(\mathbf{x})$ at the central viewpoint $(u_{0},v_{0})$ and image $L_{(u_{i},v_{j})}(\mathbf{x})$ at viewpoint $(u_{i},v_{j})$ can be described as follows:
$L_{(u_{0},v_{0})}\left(\mathbf{x}\right)=L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot d(\mathbf{x})\right),$ (2)
where $\Delta=\left(u_{i}-u_{0},v_{j}-v_{0}\right)$ represents the distance between viewpoints $(u_{i},v_{j})$ and $(u_{0},v_{0})$ in the $u$ and $v$ directions.
Certain factors such as occlusion and camera noise render Eq.(2) invalid for real-world scenarios. Therefore, this study introduces the propagation mask $m_{(u,v)}(\mathbf{x})$, with $m_{(u,v)}(\mathbf{x})=0$ in cases of occlusion or noise, and $m_{(u,v)}(\mathbf{x})=1$ otherwise. Based on this, the relationship between the central view image $L_{(u_{0},v_{0})}\left(\mathbf{x}\right)$ and the view image $L_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ in the real scenes is adjusted as follows:
$\begin{split}L_{(u_{0},v_{0})}(\mathbf{x})=m_{(u_{i},v_{j})}(\mathbf{x})\cdot L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot d(\mathbf{x})\right).\end{split}$ (3)
III-B1 Forward problem modeling
The light field is regularly sampled on the viewpoint plane, ensuring that the pixels in the central view image are visible in at least half of the viewpoints, resulting in a stronger coupling between the central view and other views. Therefore, this study employs the disparity propagation process from other views to the central view in the light field to model disparity reconstruction. This process effectively utilizes the constraints of regular sampling on the viewpoint plane, while preserving the characteristics of the light field. This provides a robust forward problem model for disparity reconstruction.
To account for occlusion and camera noise in real scenes with a light field, a forward problem model for disparity reconstruction under angular consistency was established as follows:
$L(d(\mathbf{x}))=\frac{\sum_{(u_{i},v_{j})\in\phi}\left(m_{(u_{i},v_{j})}(\mathbf{x})\cdot L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot d(\mathbf{x})\right)\right)}{\sum_{(u_{i},v_{j})\in\phi}m_{(u_{i},v_{j})}(\mathbf{x})}.$ (4)
where $\phi$ is the set of viewpoints.
III-B2 Inverse problem solving
Loss functions for selected viewpoints.
Traditional stereo-matching algorithms often utilize region-based matching windows instead of individual pixels to enhance the robustness of the matching results[39]. To exploit the local structural information of the light field fully, this study utilized a combination of L1 and mean structural similarity (MSSIM) losses [40]. Total variation (TV) [41] regularization is introduced as a loss term to exploit the sparsity of disparities in the gradient domain.
Given the light field $L_{(u,v)}(\mathbf{x})$ and the coordinate $\mathbf{x}$ of the central view, this study aims to determine the optimal parameters $\Theta^{*}$ by minimizing the objective function:
$\begin{split}&{\Theta^{*}}=\underset{\Theta}{\operatorname{arg}\operatorname{min}}\left|L_{(u_{0},v_{0})}\left(\mathbf{x}\right)-L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)\right|+\\ &\alpha\cdot\left(1-\text{MSSIM}\left(L_{(u_{0},v_{0})}\left(\mathbf{x}\right),L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)\right)\right)+\beta\cdot\Psi\left({\hat{d}_{\Theta}\left(\mathbf{x}\right)}\right),\end{split}$ (5)
where $\hat{d}_{\Theta}\left(\mathbf{x}\right)$ represents the disparity predicted at point $\mathbf{x}$ by the network and $L\left(\hat{d}_{\Theta}\left(\mathbf{x}\right)\right)$ is computed value using Eq.(4). $\Psi(\mathbf{x})$ denotes the regularization term for the parameterized disparities, where TV regularization is employed. $\alpha$ and $\beta$ are the two hyperparameters.
This study utilized the L1 + $\alpha\cdot\left(\text{MSSIM}\right)$ metric to evaluate the disparity between viewpoints that inadequately convey information and the central view image. This metric, denoted as distance $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$, is defined as follows:
$\begin{split}&E_{(u_{i},v_{j})}\left(\mathbf{x}\right)=\left|L_{(u_{0},v_{0})}\left(\mathbf{x}\right)-L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot\hat{d}_{\Theta}(\mathbf{x})\right)\right|\\ &+\alpha\cdot\left(1-\text{MSSIM}\left(L_{(u_{0},v_{0})}\left(\mathbf{x}\right),L_{(u_{i},v_{j})}\left(\mathbf{x}+\Delta\cdot\hat{d}_{\Theta}(\mathbf{x})\right)\right)\right).\end{split}$ (6)
Given that occluded pixels can effectively transmit information from at least half of the viewpoints, this study proposes a simplified strategy for computing an occlusion mask during disparity propagation. Assuming that all viewpoint masks hold a value of 1, signifying that all viewpoints contribute valuable information, this study proceeds by selecting half of the viewpoints based on a comparison of the ray distances $E_{(u_{i},v_{j})}(\mathbf{x})$. It then selects viewpoints $\Phi_{\text{selected}}$ from which the ray distance is smaller than that of the remaining viewpoints; these selected viewpoints are subsequently utilized for conducting backward propagation of rays. In general, this loss function is expressed by
$\mathcal{L}=\sum_{(u_{i},v_{j})\in\Phi_{\text{selected}}}\left(E_{(u_{i},v_{j})}\left(\mathbf{x}\right)\right)+\Psi\left({\hat{d}_{\Theta}\left(\mathbf{x}\right)}\right).$ (7)
This approach effectively avoids the complex computation of propagation masks during disparity propagation, while simultaneously reducing interference from occlusion and noise factors. Therefore, we term this approach the viewpoint selection loss strategy, which improves the accuracy of depth estimation in occluded scenes by giving precedence to the efficacy of information transmission.
Network architecture
To enhance the computational efficiency and expressive power of neural networks, recent research [42, 43, 28, 29, 30] has explored the possibility of the structured storage of latent features. While fully connected networks combined with positional encoding can record data information through weights, they reduce computational efficiency and limit the network’s capacity to represent complex patterns. Previous studies explored different approaches for storing latent features, such as 3D grids [42]. However, this grid structure suffers from resource wastage because a large portion of the internal units remain empty, resulting in the effective utilization of only surface information. To address this issue and reduce memory and training data requirements, prior work has explored the use of tree-like data structures similar to those of Acorn [43]; however, this approach incurs a larger overhead in dynamically maintaining the tree structure during training. Thus, to overcome these limitations, [28] proposed the use of multi-resolution grids for feature storage.
Considering the disparity characteristics, adjacent pixels often have similar values, reflecting the smoothness of scene depth variations. This necessitates the precise capture of subtle changes using the NDF model. Additionally, the distribution of gradients in the disparity exhibits sparsity, with most pixels having gradients close to zero, which is significant only at the scene or object edges. This sparsity implies that the local pixels share the same features, obviating the need to learn unique feature descriptions for each independent pixel, thereby reducing the training time. Thus, this study constructed an NDF network inspired by the approach proposed in [28]. Its structure comprises a multiresolution grid feature module and a MLPs module, as depicted in Fig. 2. The multiresolution grid feature module, a trainable component, encompasses $L$ grids with resolutions linearly increasing from 32 to 128, and is utilized for feature vector storage. The MLPs module consisted of fully connected layers and LeakyReLU activation functions, comprising two hidden layers, each with 256 neurons.
Figure 2: Neural-disparity-field network architecture. The network architecture comprises multiresolution grid-feature and MLPs modules. The multiresolution grid-feature module is a trainable module that stores Hash encoding feature vectors. MLPs module comprises fully connected layers and LeakyReLU activation functions.
To train the NDF, the following steps are executed:
  1. 1.
    A set of 2D sampling points $\mathbf{x}$ was extracted from the coordinate system of the reference view.
  2. 2.
    The multiresolution grid feature module was utilized to extract the corresponding features of the sampling points. These features were input into the MLPs module to predict the corresponding disparity $\hat{d}_{\Theta}\left(\mathbf{x}\right)$.
  3. 3.
    The distance $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ is calculated using Eq.(6) and retain those with $E_{(u_{i},v_{j})}\left(\mathbf{x}\right)$ smaller than the corresponding rays of the other half of the viewpoints.
  4. 4.
    The loss of the retained rays is computed using Eq.(7). Subsequently, backward propagation is performed to endow the network with the ability to learn the disparity.
IV Experiments
To comprehensively assess the performance of the proposed method, we conducted comparative experiments on several light field datasets, including the HCI 4D light field dataset [44], 4D Light Field Benchmark (CVIA Konstanz & HCI Heidelberg) [45], and Stanford Lytro dataset [46].
We evaluated the proposed method in synthetic environments of the HCI 4D light field dataset and 4D Light Field Benchmark to verify its fundamental performance and generalization. We then applied it to real Lytro-captured light-field data for a practical disparity estimation assessment. The ability of this method to directly reconstruct disparities from a network at diverse spatial resolutions was also demonstrated. Finally, an ablation study investigated the impact of combined losses, offering deep theoretical insights and practical guidance for algorithm understanding and optimization.
Experiments were conducted on hardware featuring an NVIDIA GeForce RTX 3080 with 10GB memory. The experimental parameter settings in this section are as follows: a grid feature module with six resolutions linearly increasing within the range from 32 to 128, an MSSIM loss window size of 11 pixels, and hyperparameter values $\alpha$ and $\beta$ ranging between $1\text{e}^{-1}$ and $1\text{e}^{1}$, with slight variations for different scenes. During network training, Gaussian noise with intensities ranging from $1\text{e}^{-2}$ to $1\text{e}^{0}$ was introduced.
IV-A Algorithm analysis and evaluation using synthetic data experiments
IV-A1 Complex scene experiments
In this section, disparity estimation is conducted using synthetic light-field datasets, and compared with the MV [47], LF_OCC [10], SPO [12], and AMN [14] algorithms for evaluation. Experimental results are illustrated in Fig. 3, which shows four scenes from the two datasets.
Figure 3: Disparity results for complex synthetic scenes, with odd rows showing estimated disparity maps and even rows displaying profile curves (red line in reference view) comparing actual (red) and algorithm-estimated (blue) disparities. Scenes include from HCI [44] (Buddha and Papillon) and 4D Light Field Benchmark [45] (Pens and Tomb).
In the Buddha scene, the prominent elements include a Buddha statue, slanted wooden planks, stone pillars, and dice. The intricate textures of Buddha statues and stone pillars pose challenges to the robustness of these algorithms. However, as shown in Fig. 3, all of the aforementioned algorithms effectively estimated the disparities of these elements. Slanted wooden planks tested the ability of the algorithms to handle inclined surfaces, where the proposed method outperformed MV, producing smoother disparity predictions on slanted surfaces. Additionally, the black smooth regions on the dice present challenges to the algorithm performance, where the proposed method demonstrates notable advantages over SPO in handling textureless smooth areas. The profile graphs in the second row of Fig. 3 show that the proposed method exhibited minimal oscillations, except for slight oscillations at the boundaries. In the Papillon scene, the primary elements are butterflies and leaves. The butterfly tail features a weakly textured black area and handling such a region is difficult. The proposed method successfully computed disparities in this area compared with MV, SPO, and AMN. Furthermore, for the boundary areas of the leaves, LF_OCC generates considerable noise owing to occlusion, whereas the proposed method exhibits advantages in handling such occluded regions. The Pens scene comprised multiple Pens arranged at various angles and depths. LF_OCC introduces substantial noise at pen boundaries, whereas the proposed method and other comparative algorithms effectively reconstruct the disparities in this scene. The Tomb scene contains several stone statues, and the results reconstructed using the proposed method and SPO exhibit smoothness closer to the ground truth.
This section provides detailed analyses of the Buddha, Papillon, Pens, and Tomb scenes, covering common scenarios of inclined surfaces, smooth areas, occluded regions, and texture structures. Through comparative analysis, the proposed method demonstrated advantages in handling smooth and weakly textured areas. From the profiles, it is evident that the proposed method maintains minimal deviation across the entire curve, although it exhibits oscillations, particularly at the edges of the scene.
IV-A2 Specific scene experiments
To gain a comprehensive understanding of the performance of the proposed method, we set up isolated challenges in specific scenarios: backgammons, dots, pyramids, and stripes. The Backgammon scene evaluated the algorithm’s capability in handling fine structures and occlusion boundaries. The Dots scene investigated the impact of different levels of camera noise on the reconstruction of objects of various sizes. The Pyramids scene tested the algorithm’s performance on convex, concave, circular, and planar geometrical structures, and the Stripes scene specifically assessed the influence of texture and contrast on occlusion boundary recognition. We conducted comparative experiments using the proposed method, MV, LF_OCC, SPO, AMN, EPI_ORM[20], LLFNet[22], and FastLFnet[23]. The results are presented in Fig. 4.
Figure 4: Visual comparison of disparity estimation methods across different scenes. From top to bottom, rows correspond to Backgammon, Dots, Pyramids, and Stripes scene. For each scene, the top row shows the estimated disparity, and the bottom row shows the corresponding MSE. Results (b)-(h) are derived from original data in [47].
TABLE I: Qualitative comparisons of results on synthetic datasets, presented with BP01(BadPix0.01), BP03(BadPix0.03), BP07(BadPix0.07) and MSE ($\times$ 100) metrics.
MV LF_OCC SPO AMN EPI_ORM LLFNet FastLFNet Ours
Backgammon BP01 65.842 88.839 49.940 43.599 34.316 65.975 39.839 54.878
BP03 18.547 39.826 8.639 13.354 7.238 34.337 11.409 16.378
BP07 8.924 19.006 3.782 6.913 3.988 14.853 5.138 6.656
MSE 13.225 21.587 4.587 12.287 3.411 8.058 3.986 8.929
Dots BP01 83.519 65.762 58.079 21.658 65.707 67.417 68.146 4.538
BP03 59.654 23.552 35.068 18.428 47.932 41.971 41.109 4.485
BP07 45.587 5.822 16.274 18.428 36.098 20.134 21.169 4.380
MSE 7.262 3.301 5.238 2.936 14.480 6.159 3.407 5.433
Pyramids BP01 85.392 74.144 79.206 32.485 19.063 49.817 22.188 32.675
BP03 9.824 20.495 6.263 5.643 1.301 6.495 2.193 2.853
BP07 0.776 3.172 0.861 0.449 0.324 1.059 0.620 0.129
MSE 0.051 0.098 0.043 0.023 0.016 0.035 0.018 0.016
Stripes BP01 76.372 53.669 21.879 13.499 55.135 79.080 63.400 8.389
BP03 62.417 21.257 15.460 7.966 13.943 47.050 32.594 8.053
BP07 46.192 18.408 14.987 5.835 6.871 23.837 9.442 7.518
MSE 12.173 8.131 6.955 2.128 1.744 2.390 0.984 2.518
Average BP01 77.781 70.604 52.276 27.810 43.556 65.572 48.393 25.12
BP03 37.611 26.283 16.358 11.348 17.604 32.463 21.826 7.942
BP07 25.370 11.602 8.976 7.906 11.820 14.971 9.092 4.671
MSE 8.178 8.279 4.206 4.344 4.913 4.161 2.099 4.224
Paremeteres (M) 9.1 1.8 1.4 0.5
Figures 4 (b)-(h) correspond to the different algorithms. From the experimental results, it is evident that the proposed method successfully reconstructs disparities in various challenging scenarios. In the Backgammon scene, the proposed method performed remarkably well in handling complex textures, tilted surfaces, and textureless smooth regions, particularly when compared with MV, LF_OCC, and SPO, although it failed to accurately estimate disparities at the upper-left crevice. In the Dots scene, the proposed method reconstructs more foreground points in the upper-left corner compared to the lower-right corner, indicating a certain robustness to noise; however, as the noise increases, the proposed method loses its advantage. In the Pyramids scene, all algorithms accurately reconstructed the convex and concave surfaces. The proposed method demonstrated some advantages in Stripe scene with lower demands for texture contrast. Table I presents the quantitative analyses of the scenes in Fig. 4 based on the BadPix(BP) and MSE metrics, along with the size of the network model parameters. Across various scenarios, the proposed method exhibited competitive advantages based on average quantitative metrics. Regarding the Average BP01, BP03, BP07 metrics, the proposed method shows lower misjudgment rates. Regarding the Network parameters metric, the proposed method is particularly efficient in lightweight design, having only 0.5M parameters, which makes it more suitable for resource-constrained environments than methods such as LLFNet and FastLFnet.
Although the proposed method maintains relatively low levels of BP and MSE metrics in specific scenes, such as Stripes, there is still room for further optimization, especially in handling complex textures or special structures. The proposed method demonstrates stable performance across different scenes, making it applicable to various scenarios.
IV-B Algorithm analysis and evaluation using real data experiments
Real light field data typically contain more noise, non-Lambertian radiance data, and more complex occlusion scenarios. The Stanford Lytro light field dataset [46] was captured using a Lytro Illum camera. In this study, the proposed method was compared with the built-in algorithms of the Lytro camera, MV, LF_OCC, SPO, and AMN, as shown in Fig. 5.
Figure 5: Disparity estimation results for data acquired using the Lytro Illum, with zoomed-in regions shown in close-up mode.
In Fig. 5, the first two rows depict scenes with intricate occlusion structures suitable for evaluating the performance of the algorithm in handling complex scenarios. The proposed and other methods accurately predicted the intricate branches in the foreground. LF_OCC exhibited significant errors in the background region, SPO showed multiple protrusion errors at the boundaries of the branches, and the AMN and MV algorithms were smoothed out in detail, which can be observed in the magnified area indicated by the red box in the second row. In Fig. 5, the foreground in the last two rows is mainly comprised of iron fences. Apart from the built-in algorithm of the Lytro camera, both the proposed and other algorithms accurately reconstructed the iron fences in the foreground. However, in the background region, the proposed algorithm predicted the least noise, which was particularly notable in the scene reconstructed in the last row.
The experimental results demonstrate that the proposed method accurately predicts the disparity structure in real-world scenes and reconstructs more details in edge-occlusion areas.
IV-C Spatial super-resolution experiments
The proposed method parameterizes the continuous disparity, which provides a unique advantage because a discrete system matrix does not limit it. This allows the extraction of high-resolution disparity images from a well-trained network, thereby satisfying the need for super-resolution disparity image reconstruction. Figure 6 illustrates the differently sized reconstructed images extracted from the trained network using the Dino scene. This indicates that the quality of the disparity images was not compromised when extracting high-resolution images from the network.
Figure 6: Experimental results on super-resolution: Using the Dino from the 4D Light Field Benchmark [45]. (a) presents the reference view of the scene, and (b) displays the ground truth. (c) exhibits disparity reconstructed at resolutions of $512\times 512$, with (d) showing its profile curves at the 300th line. Panels (e) and (f) exhibit the disparity reconstructed at resolutions of $1024\times 1024$ and $2048\times 2048$, respectively.
IV-D Ablation studies
This section validates the design and parameter selection through an ablation study of the combination of loss functions and viewpoint selection loss strategy.
The ablation study results presented in Fig. 7 and Table II provides evidence for the rationale behind the proposed method of combining loss functions and designing loss functions for viewpoint selection. Table II presents the quantitative analysis results for the scenes depicted in Fig. 7. These scenes included a grid-cloth background, a pear with a bee pattern, and a wooden sphere with a black pattern. Analysis of the sphere’s disparity in the images reveals that using only the L1 loss leads to noticeable layering and rough surfaces on continuous surfaces. However, the predictions using both L1 and MSSIM losses resulted in smoother surfaces. Furthermore, the inclusion of TV regularization, which imposes sparse constraints on the gradient domain of the disparity, leads to smoother surfaces.
Figure 7: Ablation experiments using the Dino from HCI [44]. The first row is the disparity, whereas the second row (excluding (a)) is the profile curves. The ground truth curves are marked in red in the profile plots, and those reconstructed by the proposed method are shown in blue. In column (a), the first row is the ground truth and the second row is the reference view of the scene. (b) The results using the L1 norm; (c) the results using both L1 and MSSIM losses. (d) and (e) show the results using L1 and MSSIM losses along with TV regularization. Panels (b)-(d) use only half of the viewpoints for backpropagation, whereas panel (e) utilizes all viewpoints.
TABLE II: Quantitative analysis of ablation experiment, presented with BadPix0.03, BadPix0.05, BadPix0.07 and MSE ($\times$ 100) metrics (smaller scores are better).
StillLife MSE BadPix0.07 BadPix0.05 BadPix0.03
L1 0.87 10.42 21.07 43.92
L1+MSSIM 0.56 8.58 14.87 29.93
TV 0.52 5.01 8.46 21.50
Full view 1.55 7.28 9.60 16.32
As shown in Table II, the combination of MSSIM loss and TV constraint yielded the lowest MSE and BadPixel values, indicating a higher reconstruction accuracy. Figure 7 (d) applies a loss function based on viewpoint selection, transmitting only half of the viewpoints during each backpropagation step, whereas Fig. 7 (e) utilizes all the viewpoints. The profile curves in Fig. 7 demonstrate that, when using the complete set of viewpoints for backpropagation, the occlusion caused by the wings of the bee cannot be effectively removed. According to Table II, employing all viewpoints for backpropagation (full view row) results in higher values for MSE, BadPix0.07, and BadPix0.05 metrics compared to propagating only half of the viewpoints (TV row), indicating poorer reconstruction quality. In disparity reconstruction from light fields, occlusions and noise within a scene are the main sources of error, although these occluded and noisy regions account for a relatively small fraction of the overall scene. During the backpropagation process, the use of more viewpoints when propagating unoccluded areas improves the reconstruction accuracy. Consequently, regarding the BadPix0.03 metric, employing all available viewpoints for backpropagation (full view row) yielded lower values than using only half of the viewpoints (TV row), indicating superior reconstruction quality.
V Discussion
We conducted comparative experiments on widely used light-field datasets to validate the performance of the proposed method in disparity estimation. The results demonstrate the exceptional proficiency of the proposed method in addressing inclined surfaces, smooth textureless regions, and occlusions in intricate scenes. When confronted with the weakly textured tail area of the butterfly and occluded leaf boundaries in the Papillon scene, the proposed method delivered a higher accuracy and smoother disparity estimates than MV, LF_OCC, SPO, and AMN. Further validation of the method’s resilience and adaptability was sought through tailored challenges in scenes such as backgammons, dots, pyramids, and stripes, which tested its ability to handle fine structures, noise interference, convex-concave geometries, and low-texture contrast scenarios. Despite a slight decline in the performance under high-noise conditions, the proposed method achieved effective disparity reconstruction and detail preservation in most cases. Additionally, experiments on the Stanford Lytro light field dataset demonstrated the success of the proposed method in accurately discerning disparity structures, even in highly noisy and complex occluded real-world scenes, particularly in suppressing background noise and reconstructing foreground edge details relative to contrastive algorithms. Moreover, by extracting disparities at varying resolutions directly from a pre-trained NDF, the proposed approach transcends the conventional limitations imposed by the reliance on discrete matrix-based disparity representations. Finally, the ablation studies shed light on the influence of combination strategies and selected viewpoints on the loss function. The MSSIM loss curbed the spread of local pixel errors. In contrast, including total variation regularization enforces constraints on the disparity gradient domain, ensuring smooth and coherent estimation outcomes. The capability to propagate with only half of the viewpoints bolsters the resistance to occlusion and noise.
Although the proposed method demonstrated excellent performance in the above experiments, the disparity map exhibited higher frequencies than the natural images, particularly in edge regions with disparity differences, resulting in oscillations in the predicted disparities. Thus, further research is required to investigate parameterized network models for disparities to enhance algorithm performance and applicability.
In summary, the proposed compact network model demonstrates efficiency and robustness in handling weak textures, varying noise levels, and occlusion scenarios. This could potentially provide more precise and reliable solutions for light field depth estimation in practical applications.
VI Conclusions
We proposed a novel iterative approach to reconstruct the NDF that utilized neural network parameterization to achieve an implicit, continuous disparity function. The essential feature of this method is the implicit parameterization of the disparity by neural fields, which overcomes the resolution limitations inherent in traditional light-field disparity estimation. Compared to conventional methods based on discrete disparity maps, the proposed approach significantly reduced computational complexity while maintaining high precision and demonstrating superior robustness, particularly in handling complex scenes where occlusion and noise affect accurate disparity reconstruction. The proposed method employed a differentiable framework to construct a neural network model capable of modeling disparity continuously, thereby effectively avoiding the loss of detail caused by discrete representations. During the backpropagation optimization process, only a subset of rays was selected for propagation computation in the viewpoint direction, helping to mitigate occlusion effects and noise interference in the disparity estimation. Experimental validation indicates that the reconstructed NDF achieved accurate disparity on both synthetic and real datasets, enabling the effective recovery of high-resolution disparity.